OpenStudy (anonymous):

56s^2+23s-63 Need help solving this one

OpenStudy (bahrom7893):

= 0 ?

OpenStudy (bahrom7893):

56s^2+23s-63 = 0?

OpenStudy (anonymous):

Nope have to factor this out. I know the common denominator is 7 but I get lost and confused after that .

OpenStudy (bahrom7893):

oh just use the discriminant..

OpenStudy (bahrom7893):

D = b^2 - 4ac

OpenStudy (anonymous):

(7s+9)(8x-7) if I did it right.

OpenStudy (anonymous):

Ok this is what I get 56s^2+23s-63 (56s+ )(s- )

OpenStudy (anonymous):

oops I meant for the second one to be an s as well

OpenStudy (bahrom7893):

D = 529 + 14112

OpenStudy (anonymous):

Now got to figure out what the numbers are for the remaining. I know that 9*7 is 63 But when you work it into the problem and then use the foil method I dont get the 23s in the middle Have also tried 12*3

OpenStudy (bahrom7893):

x1 = (- 23 - 121)/(2*56) = 9/7

OpenStudy (anonymous):

(7s+9)(8s-7)

OpenStudy (bahrom7893):

x2 = (- 23 + 121)/(2*56) = 7/8

OpenStudy (anonymous):

TBates you rock!!!!!

OpenStudy (anonymous):

I used the quadratic equation to find one of the zeros and worked backwards to get the factors.

OpenStudy (anonymous):

Thank you thank you thank you !

OpenStudy (bahrom7893):

well that's what i suggested.. wasn't it?

OpenStudy (bahrom7893):

(x - 9/7)(x - 7/8)

OpenStudy (anonymous):

That it is. The quadratic formula will always work!

OpenStudy (anonymous):

yes it was you were correct as well